OpenStudy (anonymous):

Simplify 3/x + 2/y over 3/x - 2/y

OpenStudy (anonymous):

\[\frac{3/x + 2/y}{3/x - 2/y} =\frac{\frac{3y+2x}{6}}{\frac{3y-2x}{6}}\] \[=\frac{3y+2x}{3y-2x}\]

OpenStudy (anonymous):

thanks :)